# MAPS: A Multilingual Benchmark for Global Agent Performance and Security

### v1

@@ 4.2 Results
From part (b), we can see that there is no clear correlation between multilingual security robustness (ASB) and multilingual performance degradation. This disconnect is especially clear in real-world, language-heavy tasks like GAIA, where performance drops sharply, while structured tasks like SWE-bench and MATH remain largely unaffected. This highlights that multilingual security alignment does not directly track with multilingual task accuracy, notably in language-rich agentic tasks.
5 Discussion
This section presents practical recommendations for multilingual agent deployment and directions for advancing the benchmark in future work.
5.1 Guidelines for Multilingual Evaluation and Risk Assessment
Language-Aware Deployment Guidelines. Before deploying an AI agent in a multilingual setting, analyze the linguistic composition of its expected input, particularly the balance between structured elements (e.g., code, formal queries) and localized natural language. Inputs with a high proportion of non-English content, especially those involving less formalized or more natural language, tend to increase the risk of performance and safety degradation. We therefore recommend that for any such case, developers should conduct a Multilingual Benchmark Assessment using a diverse, language-sensitive evaluation suite, such as ours, for AI agents operating across languages. This helps reveal hidden vulnerabilities and promotes reliable real-world behavior in multilingual conditions.
Prioritize Multilingual Adaptation by Task Type. Our findings suggest that the need for multilingual adaptation in agentic systems should be guided by task type. For structured tasks with minimal linguistic variability, such as coding, cross-lingual transfer can often be achieved with minimal adjustment. However, for complex, real-world tasks or safety-critical decisions (e.g., GAIA, ASB), multilingual robustness remains limited, and thus, dedicated multilingual alignment and adaptation are essential. MAPS offers a practical framework to identify where multilingual adaptation is needed, helping prioritize resource allocation for post-training based on task-specific language sensitivity.
Multilingual Inputs Amplify Agentic Security Vulnerabilities. Our evaluation on ASB revealed that multilingual adversarial inputs can bypass agent safety mechanisms with minimal sophistication. Direct translations of English jailbreak prompts—without any adaptation or obfuscation—were sufficient to induce policy-violating behavior in multiple languages. This highlights a critical risk: even simple adversarial prompts become significantly more effective when the input is localized, and are often sufficient to exploit security vulnerabilities in AI agents. Developers of safety-critical agentic systems should treat multilingual robustness as a core security concern and include translated prompts in safety evaluations using benchmarks like ours.
5.2 Benchmark Limitation
While MAPS represents the first multilingual suite for evaluating agentic AI systems, there are natural opportunities for future expansion. The current release includes four datasets, one agent per dataset, and ten target languages, offering a strong foundation for assessing multilingual robustness. Extending coverage to additional domains such as healthcare or legal reasoning, as well as incorporating multiple agents and extremely low-resource languages (e.g., Amharic or Uyghur), would further enhance the benchmark’s scope and relevance. Nonetheless, the current suite already surfaces clear trends in performance and security degradation across languages, offering valuable insights for guiding multilingual deployment. We view this work as a meaningful starting point and invite the community to build on our open-source release to advance more inclusive and resilient agentic AI systems.
6 Conclusions
We introduce the first multilingual benchmark suite for evaluating agentic AI systems, addressing a critical gap in assessing language-specific performance and safety limitations. By adapting and localizing four widely used agentic benchmarks—GAIA, SWE-bench, MATH, and ASB—into ten diverse languages, our suite enables the analysis of agent behavior under multilingual conditions.
Constructed through a hybrid translation pipeline and human verification, the benchmark ensures high linguistic fidelity and structural consistency. Experimental results reveal consistent degradation in both performance and robustness when agents operate in non-English settings, particularly in tasks involving natural language reasoning and safety-critical behavior.
These findings underscore the importance of language-aware evaluation and targeted multilingual adaptation, especially for real-world agentic deployments. We view this benchmark as a practical and extensible foundation for building more inclusive, resilient, and globally reliable agentic AI systems, and we invite the community to build upon it.

### v2

@@ A.6 Results Additional Information
This section supplements the main manuscript with detailed quantitative and qualitative results. While the main paper summarizes overall trends and multilingual effects, Tables 1114 provide per-language scores, averages, and standard deviations for each benchmark. Fig. 9 provides a multilingual qualitative failure example of Hindi language in GAIA, including a comparison of main steps in English and Hindi. Fig.  10 provides a language-wise breakdown of relative performance degradation across datasets. From the Figure, we can see that there is no clear correlation between multilingual security robustness (ASB) and multilingual performance degradation. This disconnect is especially clear in real-world, language-heavy tasks like GAIA, where performance drops sharply, while structured tasks like SWE-bench and MATH remain largely unaffected. This highlights that multilingual security alignment does not directly track with multilingual task accuracy, notably in language-rich agentic tasks.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.5094 0.4450 0.4695 0.4746 0.0325
Portuguese (br) 0.3902 0.3658 0.3597 0.3719 0.0161
German 0.3658 0.3475 0.3597 0.3577 0.0093
Spanish 0.3475 0.3536 0.3680 0.3564 0.0105
Italian 0.3536 0.3353 0.3170 0.3353 0.0183
Russian 0.4268 0.4085 0.3530 0.3961 0.0384
Arabic 0.2866 0.3292 0.3170 0.3109 0.0219
Hebrew 0.3818 0.3597 0.3353 0.3589 0.0233
Hindi 0.3353 0.3780 0.3353 0.3496 0.0247
Korean 0.3597 0.3231 0.3475 0.3434 0.0186
Japanese 0.4180 0.3597 0.3170 0.3649 0.0507
Table 11: GAIA benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.29 0.33 0.31 0.31 0.020
Portuguese (br) 0.32 0.29 0.25 0.323 0.0351
German 0.29 0.27 0.30 0.287 0.0153
Spanish 0.31 0.31 0.31 0.31 0.0000
Italian 0.30 0.29 0.32 0.303 0.0150
Russian 0.31 0.34 0.34 0.33 0.0173
Arabic 0.33 0.28 0.31 0.307 0.0252
Hebrew 0.31 0.27 0.27 0.283 0.0231
Hindi 0.28 0.28 0.28 0.280 0.0000
Korean 0.30 0.32 0.32 0.313 0.0115
Japanese 0.31 0.29 0.29 0.297 0.0153
Table 12: SWE-Bench benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.514 0.514 0.514 0.514 0.000
Portuguese (br) 0.514 0.514 0.514 0.514 0.000
German 0.500 0.500 0.500 0.500 0.000
Spanish 0.536 0.536 0.536 0.536 0.000
Italian 0.479 0.479 0.479 0.479 0.000
Russian 0.507 0.507 0.507 0.507 0.000
Arabic 0.514 0.514 0.514 0.514 0.000
Hebrew 0.521 0.521 0.521 0.521 0.000
Hindi 0.436 0.443 0.443 0.440 0.004
Korean 0.414 0.421 0.421 0.419 0.004
Japanese 0.443 0.443 0.443 0.443 0.000
Table 13: MATH benchmark performance across languages.
Language Run #1 Run #2 Run #3 Mean StdDev
English (baseline) 0.534 0.536 0.532 0.534 0.002
Portuguese (br) 0.648 0.657 0.654 0.653 0.005
German 0.659 0.652 0.654 0.655 0.004
Spanish 0.595 0.613 0.597 0.602 0.010
Italian 0.612 0.618 0.603 0.611 0.008
Russian 0.636 0.646 0.642 0.641 0.005
Arabic 0.515 0.504 0.511 0.510 0.005
Hebrew 0.558 0.571 0.572 0.567 0.008
Hindi 0.617 0.616 0.614 0.616 0.002
Korean 0.590 0.589 0.588 0.589 0.001
Japanese 0.484 0.489 0.497 0.490 0.007
Table 14: ASB benchmark performance across languages.